Exercise 3 (3 points): For each of the following make sure to show your calculations: a. Suppose that there are N-10 firms in a market, and each of them have an equal share of the market, sp that s.-1/N-1/10-10% for each firm. What is the HHI for this market? Characterize this market Structure based on the value of the HHI. b. Suppose instead that one firm in this market dominates such that it has 90% of the market share, while the remaining 9 firms each have an equal share of the remainder. What is the HHI in this case? Characterize this market struciure on the value of the HHI. Interpret how this result compares to case (a) above. suppose a new fim enters che maiket and pulls away market share from each of the incumbent firms such that the sharesbecome s-1/11-909%. What js the Hill in this case? Characterize this market structure based on th of thé HIHI. Interpret how this result compares to case () above. Page 2 of 2
Answer #1

Use the formula

Herfindahl-Hirschman Index is used to measure the industrial concentration which is the sum of squared market shares of all t

a) HHI = (10)^2 + (10)^2 ..... + (10)^2

= 10*100 = 1000

Since HHI is less than 1500, these firms are competitive. Markets that compete in competition have a lower value of four-firm concentration ratio. The HHI value for these markets is less than 1,500.

b) HHI = (90)^2 + (1)^2 ..... + (1)^2

= 8100 + 1 = 8101

Since this market has a higher value of HHI. it is very close to a monopoly. Monopolies have an HHI close to 10000.

c) HHI = 9.09^2 +  9.09^2 + ... +  9.09^2

= 10* 9.09^2

= 826.28

Since HHI is less than 1500, these firms are competitive. The HHI value for these markets is less than 1,500.